# Reference Grounded Skill Discovery

### v2

@@ Abstract
Scaling unsupervised skill discovery algorithms to high-DoF agents remains challenging. As dimensionality increases, the exploration space grows exponentially, while the manifold of meaningful skills remains limited. Therefore, semantic meaningfulness becomes essential to effectively guide exploration in high-dimensional spaces. In this work, we present Reference-Grounded Skill Discovery (RGSD), a novel algorithm that grounds skill discovery in a semantically meaningful latent space using reference data. RGSD first performs contrastive pretraining to embed motions on a unit hypersphere, clustering each reference trajectory into a distinct direction. This grounding enables skill discovery to simultaneously involve both imitation of reference behaviors and the discovery of semantically related diverse behaviors. On a simulated SMPL humanoid with $359$-D observations and $69$-D actions, RGSD successfully imitates skills such as walking, running, punching, and sidestepping, while also discover variations of these behaviors. In downstream locomotion tasks, RGSD leverages the discovered skills to faithfully satisfy user-specified style commands and outperforms imitation-learning baselines, which often fail to maintain the commanded style. Overall, our results suggest that lightweight reference-grounding offers a practical path to discovering semantically rich and structured skills in high-DoF systems.
@@ 1 Introduction
Figure 1: Comparison of learned skills from METRA and RGSD. Our method can discover structured skills in high-DoF systems.
@@
The ultimate goal of unsupervised skill discovery is to acquire a set of reusable skills that can be applied to arbitrary downstream tasks. Achieving this goal requires the learned skills to satisfy two key desiderata: diversity and semantic meaningfulness. First, the skill set should be sufficiently diverse to cover the wide distribution of possible downstream tasks. For example, a humanoid agent should be equipped with both turn-left and turn-right skills; if it only learns to walk-forward, it will fail on general locomotion tasks, such as goal reaching or heading. Second, the skills must be semantically meaningful, since downstream tasks of interest are typically defined in semantic terms. For instance, a humanoid might acquire a skill that produces a meaningless jittering behavior, such as vibrating both arms in the air. While this behavior may be diverse, it is difficult to imagine a task where such a skill would be useful. Therefore, it is crucial that discovered skills are not only diverse but also semantically interpretable.
@@
Recent progress in skill discovery has demonstrated the ability to learn skills that satisfy both desiderata, particularly in relatively low degree-of-freedom (DoF) systems such as benchmark environments (diayn; vic; cic), quadrupeds (cheng2024learning; atanassov2024constrained; rho2025unsupervised; cathomen2025divide), or simple robotic manipulators with grippers (park2023controllability; rholanguage). However, to the best of our knowledge, scaling unsupervised skill discovery to high-DoF agents remains an open challenge. As the DoF increases, the exploration space grows exponentially, while the portion of the semantically meaningful manifold remains relatively small. This mismatch renders skill discovery in high-DoF systems particularly difficult. As shown in Fig. 1, skills learned by the state-of-the-art unsupervised skill discovery algorithm, METRA (metra), fail to yield meaningful behaviors. Joints move randomly, producing highly unstructured motions in which arms, legs, torso, and head move independently and arbitrarily. This highlights the necessity of a mechanism that grounds skill discovery within a meaningful manifold.
@@
Our key insight is that to tame the curse of dimensionality in high-DoF skill discovery, we need to construct a semantically meaningful skill latent space a priori and constrain exploration within this space. In this paper, we propose Reference-Grounded Skill Discovery (RGSD), a simple yet effective algorithm that achieves this by leveraging reference data to ground the latent skill space before exploration begins. Building upon DIAYN (diayn), RGSD first uses contrastive learning on reference trajectories to embed motions on a unit hypersphere, where each reference behavior corresponds to a distinct direction. This grounding step establishes a semantically meaningful manifold that constrains subsequent exploration. RGSD then utilizes this pre-structured latent space to simultaneously imitate reference skills and discover novel behaviors that are semantically related to the references. This two-stage approach is analogous to recent large language model training regimes (guo2025deepseek; yang2025qwen3), where pretraining with self-supervised objectives establishes a meaningful exploration space before reinforcement learning fine-tuning. As a result, RGSD not only reproduces motions present in the reference data but also discovers new skills that emerge as coherent variations of existing ones, rather than arbitrary high-dimensional movements.
@@ 2.1 Unsupervised skill discovery
The most widely adopted approach for acquiring diverse skills is based on maximizing the mutual information (MI) $\mathcal{I}(S;Z)$ between a latent variable $z$ and the agent’s states $s$ (vic; diayn; hansen2019fast; liu2021aps; cic). However, because MI is typically estimated via KL divergence, it suffers from a key limitation: even minor differences between skills can fully maximize the objective, as long as the discriminator can distinguish them. To address this, methods based on the Wasserstein Dependency Measure (WDM, ozair2019wasserstein) have been proposed (he2022wasserstein; lsd; metra), which explicitly seek maximal differences between skills. Nevertheless, our method builds on MI-based approaches, and in Section 6 we explain why extending WDM-based algorithms with our idea faces inherent challenges. Another relevant line of work is CIC (cic), which, like our approach, employs contrastive learning. The key difference is that CIC optimizes a contrastive objective on online data gathered by the agent, leaving it with the same scaling difficulties in high-DoF systems.
More recently, approaches such as LGSD (rholanguage) and DoDont (kim2024s) have explored injecting guidance from LLMs or videos to semantically ground skill discovery. While motivated by a similar goal, these methods still struggle to scale to high-DoF systems, as their guidance is provided only at a high level. In contrast, our method leverages reference data directly, enabling a more effective grounding mechanism that scales to complex, high-DoF agents.
@@ 2.2 Imitation-based skill acquisition
Imitation learning offers another major paradigm for acquiring a repertoire of skills. A common approach is based on GAIL (ho2016generative), which uses adversarial training to align the marginal state distribution of the learned policy with that of expert demonstrations. Building on this idea, ASE (peng2022ase) learns a controllable skill embedding from large-scale human motion data using a GAIL-style objective. CALM (tessler2023calm) mitigates the mode collapse observed in ASE by incorporating a conditional discriminator and a motion encoder, enabling skill generation conditioned on specific reference motions. More recently, Meta-Motivo (tirinzoni2025zero) combines GAIL with forward–backward representation learning (blier2021learning; touati2021learning) to extract a skill set tailored for zero-shot RL.
@@ 3 Preliminaries
A common formulation is to maximize the mutual information (MI) $\mathcal{I}(S;Z)$ between skills and visited states. Prior work (vic; diayn; cic; liu2021behavior) leverages the decomposition $\mathcal{I}(s;z)=\mathcal{H}(s)-\mathcal{H}(s\mid z)=\mathcal{H}(z)-\mathcal{H}(z\mid s).$ Because $\mathcal{H}(z\mid s)$ is intractable, diayn introduced a neural encoder $q_{\phi}(z|s)$ to obtain a variational lower bound $\mathcal{G}(\theta,\phi)$ of the skill discovery objective $\mathcal{F}(\theta)$:
@@ 4.1 Pretraining: grounding latent space on reference motions
and optimize the InfoNCE loss (oord2018representation):
@@ 4.2 Imitation of reference skills
This formulation can be viewed as feature-based imitation with structured representations, as proposed by li2025feature. Unlike DeepMimic (peng2018deepmimic) or MaskedMimic (tessler2024maskedmimic), which reward joint-level similarity, our method evaluates similarity in a learned latent space.
@@ 5 Experimental results and discussion
We compare RGSD against both unsupervised skill discovery (USD) methods and imitation learning (IL) based skill acquisition methods. To assess the impact of reference guidance, we include pure USD baselines. Specifically, we compare against DIAYN (diayn) and METRA (metra). DIAYN, which forms the backbone of RGSD, also serves as an ablation without reference grounding, while METRA represents a distance-maximization approach to USD that explicitly seeks maximal differences between skills.
For the IL baselines, we consider ASE, CALM, and Meta-Motivo, all of which utilize a GAIL-style objective (ho2016generative) for imitation. ASE builds on InfoGAIL (li2017infogail), combining mutual-information maximization with an adversarial imitation reward, making it the closest baseline that jointly addresses skill discovery and imitation. CALM, a variant of ASE, further augments the imitation objective using a motion-conditioned discriminator. Meta-Motivo integrates imitation rewards with forward–backward representation learning.
@@
We selected 20 reference motions from the osu_accad dataset. These motions are categorized into five tasks: walk, run, sidestep, backward, and punch, with each task containing 2 to 6 relevant motions. A complete list of motions is provided in Appendix G.3. Training was conducted in the GPU-based simulator Isaac Gym (makoviychuk2021isaac), using PPO (ppo) as the RL algorithm. The full set of hyperparameters is provided in Appendix G.4. To ensure stable training, all methods employed an early termination condition: whenever the robot fell, the episode was terminated.
@@ 5.1 Evaluation of Imitation
As shown in Table 1, RGSD achieves low Cartesian error while Meta-Motivo achieves low FID scores across most motions. In contrast, pure USD baselines fail to discover skills that closely resemble the reference dataset. While DIAYN has demonstrated the ability to discover basic locomotion behaviors in low DoF ranging from 3 to 6 environments such as HalfCheetah, Ant, and Hopper (brockman2016openai), they struggle with the 69-DoF SMPL agent, producing behaviors that are not semantically meaningful. This emphasizes the impact of reference guidance.
@@ 5.3 Downstream task evaluation
To encourage both goal reaching and adherence to the commanded style, we adopt a reward similar to tessler2023calm:
@@ 6 Understanding the challenges of expanding RGSD with METRA
However, this formulation exposes a critical issue when dealing with repetitive motions. Consider walking: the agent begins at some pose, takes a step, and eventually returns to a pose nearly identical to the starting one. Because all observations are computed in the local frame, the initial state $s_{0}$ and final state $s_{T}$ become identical. In this case, the METRA reward $(\phi(s_{T})-\phi(s_{0}))^{\top}z$ collapses to zero. As a result, repetitive behaviors cannot be framed as reward maximization under METRA, and more broadly, WDM-based approaches defined in a local coordinate system are not straightforward to capture repetitive dynamics.
@@ 7 Conclusion
  • The full codebase will be publicly released upon acceptance.
@@ Appendix A Qualitative results for controlling diversity experiment
Figure 6: Red dots indicate the agent’s trajectory. RGSD enables test-time control of behavioral diversity by varying the sampling distribution of the latent vector $z$.
@@ G.1 State and Action space
We use a 3D SMPL humanoid character, commonly adopted in prior work (Luo2023PerpetualHC; tessler2024maskedmimic; tirinzoni2025zero). The state representation is defined as follows:
@@ G.2 Implementation Details of Baseline algorithms
Finally, following conventions in skill discovery frameworks (vic; diayn; cic; metra; rholanguage), we sampled a skill latent at the beginning of each episode and kept it fixed throughout the rollout. While transitioning between skills is orthogonal and can be incorporated into both our method and the baselines to further improve robustness, we excluded it here to focus on a clean understanding of algorithmic contributions.
@@ G.4 Hyperparameters
Optimizer Adam(kingma2014adam)

### v3

@@ Abstract
Scaling unsupervised skill discovery algorithms to high-DoF agents remains challenging. As dimensionality increases, the exploration space grows exponentially, while the manifold of meaningful skills remains limited. Therefore, semantic meaningfulness becomes essential to effectively guide exploration in high-dimensional spaces. In this work, we present Reference-Grounded Skill Discovery (RGSD) [^1] Videos at seungeunrho.github.io/projects/RGSD, a novel algorithm that grounds skill discovery in a semantically meaningful latent space using reference data. RGSD first performs contrastive pretraining to embed motions on a unit hypersphere, clustering each reference trajectory into a distinct direction. This grounding enables skill discovery to simultaneously involve both imitation of reference behaviors and the discovery of semantically related diverse behaviors. On a simulated SMPL humanoid with $359$-D observations and $69$-D actions, RGSD successfully imitates skills such as walking, running, punching, and sidestepping, while also discover variations of these behaviors. In downstream locomotion tasks, RGSD leverages the discovered skills to faithfully satisfy user-specified style commands and outperforms imitation-learning baselines, which often fail to maintain the commanded style.
@@ 1 Introduction
Figure 1: Comparison of skills learned by METRA and RGSD. To our knowledge, this is the first work to discover structured skills in high-DoF systems by grounding them in reference motions.
@@
The ultimate goal of unsupervised skill discovery is to acquire a set of reusable skills that can be applied to arbitrary downstream tasks. Achieving this goal requires the learned skills to satisfy two key desiderata: diversity and semantic meaningfulness. First, the skill set should be sufficiently diverse to cover the wide distribution of possible downstream tasks. For example, a humanoid agent should be equipped with both turn-left and turn-right skills; if it only learns to walk-forward, it will fail on general locomotion tasks, such as goal reaching or heading. Second, the discovered skills must be semantically meaningful, as downstream tasks are typically defined in semantic terms. For example, consider a set of unstructured humanoid behaviors such as independent vibrating or jittering of individual body parts. Although these behaviors may be diverse, it is difficult to envision realistic tasks for which such skills would be useful. Therefore, it is essential that discovered skills are not only diverse, but also structured and semantically interpretable.
@@
Recent progress in skill discovery has demonstrated the ability to learn skills that satisfy both desiderata, particularly in relatively low degree-of-freedom (DoF) systems such as benchmark environments (Eysenbach et al., 2018; Gregor et al., 2016; Laskin et al., 2022), quadrupeds (Cheng et al., 2024; Atanassov et al., 2024; Rho et al., 2025; Cathomen et al., 2025), or simple robotic manipulators with grippers (Park et al., 2023a; Rho et al., 2024). However, scaling unsupervised skill discovery to high-DoF agents remains an open challenge. As the DoF increases, the exploration space grows exponentially, while the portion of the semantically meaningful manifold remains relatively small. This mismatch renders skill discovery in high-DoF systems particularly difficult. As shown in Fig. 1, skills learned by the state-of-the-art unsupervised skill discovery algorithm, METRA (Park et al., 2023b), fail to yield meaningful behaviors. Joints move randomly, producing highly unstructured motions in which arms, legs, torso, and head move independently and arbitrarily. This highlights the necessity of a mechanism that grounds skill discovery within a meaningful manifold.
@@
Our key insight is that to tame the curse of dimensionality in high-DoF skill discovery, we need to construct a semantically meaningful skill latent space a priori and constrain exploration within this space. In this paper, we propose Reference-Grounded Skill Discovery (RGSD), a simple yet effective algorithm that achieves this by leveraging reference data to ground the latent skill space before exploration begins. Building upon DIAYN (Eysenbach et al., 2018), RGSD first uses contrastive learning on reference trajectories to embed motions on a unit hypersphere, where each reference behavior corresponds to a distinct direction. This grounding step establishes a semantically meaningful manifold that constrains subsequent exploration. RGSD then utilizes this pre-structured latent space to simultaneously imitate reference skills and discover novel behaviors that are semantically related to the references. This two-stage approach is analogous to recent large language model training regimes (Guo et al., 2025; Yang et al., 2025), where pretraining with self-supervised objectives establishes a meaningful exploration space before reinforcement learning fine-tuning. As a result, RGSD not only reproduces motions present in the reference data but also discovers new skills that emerge as coherent variations of existing ones, rather than arbitrary high-dimensional movements.
@@ 2.1 Unsupervised skill discovery
The most widely adopted approach for acquiring diverse skills is based on maximizing the mutual information (MI) $\mathcal{I}(S;Z)$ between a latent variable $z$ and the agent’s states $s$ (Gregor et al., 2016; Eysenbach et al., 2018; Hansen et al., 2019; Liu and Abbeel, 2021a; Laskin et al., 2022). However, because MI is typically estimated via KL divergence, it suffers from a key limitation: even minor differences between skills can fully maximize the objective, as long as the discriminator can distinguish them. To address this, methods based on the Wasserstein Dependency Measure (WDM, Ozair et al. (2019)) have been proposed (He et al., 2022; Park et al., 2021; 2023b), which explicitly seek maximal differences between skills. Nevertheless, our method builds on MI-based approaches, and in Section 6 we explain why extending WDM-based algorithms with our idea faces inherent challenges. Another relevant line of work is CIC (Laskin et al., 2022), which, like our approach, employs contrastive learning. The key difference is that CIC optimizes a contrastive objective on online data gathered by the agent, leaving it with the same scaling difficulties in high-DoF systems.
More recently, approaches such as LGSD (Rho et al., 2024) and DoDont (Kim et al., 2024) have explored injecting guidance from LLMs or videos to semantically ground skill discovery. While motivated by a similar goal, these methods still struggle to scale to high-DoF systems, as their guidance is provided only at a high level. In contrast, our method leverages reference data directly, enabling a more effective grounding mechanism that scales to complex, high-DoF agents.
@@ 2.2 Imitation-based skill acquisition
Imitation learning offers another major paradigm for acquiring a repertoire of skills. A common approach is based on GAIL (Ho and Ermon, 2016), which uses adversarial training to align the marginal state distribution of the learned policy with that of expert demonstrations. Building on this idea, ASE (Peng et al., 2022) learns a controllable skill embedding from large-scale human motion data using a GAIL-style objective. CALM (Tessler et al., 2023) mitigates the mode collapse observed in ASE by incorporating a conditional discriminator and a motion encoder, enabling skill generation conditioned on specific reference motions. More recently, Meta-Motivo (Tirinzoni et al., 2025) combines GAIL with forward–backward representation learning (Blier et al., 2021; Touati and Ollivier, 2021) to extract a skill set tailored for zero-shot RL.
@@ 3 Preliminaries
A common formulation is to maximize the mutual information (MI) $\mathcal{I}(S;Z)$ between skills and visited states. Prior work (Gregor et al., 2016; Eysenbach et al., 2018; Laskin et al., 2022; Liu and Abbeel, 2021b) leverages the decomposition $\mathcal{I}(s;z)=\mathcal{H}(s)-\mathcal{H}(s\mid z)=\mathcal{H}(z)-\mathcal{H}(z\mid s).$ Because $\mathcal{H}(z\mid s)$ is intractable, Eysenbach et al. (2018) introduced a neural encoder $q_{\phi}(z|s)$ to obtain a variational lower bound $\mathcal{G}(\theta,\phi)$ of the skill discovery objective $\mathcal{F}(\theta)$:
@@ 4.1 Pretraining: grounding latent space on reference motions
and optimize the InfoNCE loss (Oord et al., 2018):
@@ 4.2 Imitation of reference skills
This formulation can be viewed as feature-based imitation with structured representations, as proposed by Li et al. (2025). Unlike DeepMimic (Peng et al., 2018) or MaskedMimic (Tessler et al., 2024), which reward joint-level similarity, our method evaluates similarity in a learned latent space.
@@ 5 Experimental results and discussion
We compare RGSD against both unsupervised skill discovery (USD) methods and imitation learning (IL) based skill acquisition methods. To assess the impact of reference guidance, we include pure USD baselines. Specifically, we compare against DIAYN (Eysenbach et al., 2018) and METRA (Park et al., 2023b). DIAYN, which forms the backbone of RGSD, also serves as an ablation without reference grounding, while METRA represents a distance-maximization approach to USD that explicitly seeks maximal differences between skills.
For the IL baselines, we consider ASE, CALM, and Meta-Motivo, all of which utilize a GAIL-style objective (Ho and Ermon, 2016) for imitation. ASE builds on InfoGAIL (Li et al., 2017), combining mutual-information maximization with an adversarial imitation reward, making it the closest baseline that jointly addresses skill discovery and imitation. CALM, a variant of ASE, further augments the imitation objective using a motion-conditioned discriminator. Meta-Motivo integrates imitation rewards with forward–backward representation learning.
@@
We selected 20 reference motions from the ACCAD dataset. These motions are categorized into five tasks: walk, run, sidestep, backward, and punch, with each task containing 2 to 6 relevant motions. A complete list of motions is provided in Appendix G.3. Training was conducted in the GPU-based simulator Isaac Gym (Makoviychuk et al., 2021), using PPO (Schulman et al., 2017) as the RL algorithm. The full set of hyperparameters is provided in Appendix G.4. To ensure stable training, all methods employed an early termination condition: whenever the robot fell, the episode was terminated.
@@ 5.1 Evaluation of Imitation
As shown in Table 1, RGSD achieves low Cartesian error while Meta-Motivo achieves low FID scores across most motions. In contrast, pure USD baselines fail to discover skills that closely resemble the reference dataset. While DIAYN has demonstrated the ability to discover basic locomotion behaviors in low DoF ranging from 3 to 6 environments such as HalfCheetah, Ant, and Hopper (Brockman et al., 2016), they struggle with the 69-DoF SMPL agent, producing behaviors that are not semantically meaningful. This emphasizes the impact of reference guidance.
@@ 5.3 Downstream task evaluation
To encourage both goal reaching and adherence to the commanded style, we adopt a reward similar to Tessler et al. (2023):
@@ 6 Understanding the challenges of expanding RGSD with METRA
However, this formulation exposes a critical issue when dealing with repetitive motions. Consider walking: the agent begins at some pose, takes a step, and eventually returns to a pose nearly identical to the starting one. Because all observations are computed in the local frame, the initial state $s_{0}$ and final state $s_{T}$ become identical. In this case, the METRA reward $(\phi(s_{T})-\phi(s_{0}))^{\top}z$ collapses to zero. As a result, repetitive behaviors cannot be framed as reward maximization under METRA, and more broadly, WDM-based approaches defined in a local coordinate system are not straightforward to capture repetitive dynamics. This observation is consistent with the concurrent findings (Park et al., ).
@@ 7 Conclusion
Acknowledgments
This research has been funded by the Industrial Technology Innovation Program (P0028404, development of a product level humanoid mobile robot for medical assistance equipped with bidirectional customizable human-robot interaction, autonomous semantic navigation, and dual-arm complex manipulation capabilities using large-scale artificial intelligence models) of the Ministry of Industry, Trade and Energy of Korea. In addition, we thank Jeonghwan Kim for the discussion.
@@ Appendix A Qualitative results for controlling diversity experiment
Figure 6: Red dots indicate the agent’s trajectory. From left to right, as we sample the latent variable $z$ from increasingly wider distributions centered around the reference motion, the resulting behaviors become more diverse while still preserving the core characteristics of the original motion.
@@ G.1 State and Action space
We use a 3D SMPL humanoid character, commonly adopted in prior work (Luo et al., 2023; Tessler et al., 2024; Tirinzoni et al., 2025). The state representation is defined as follows:
@@ G.2 Implementation Details of Baseline algorithms
Finally, following conventions in skill discovery frameworks (Gregor et al., 2016; Eysenbach et al., 2018; Laskin et al., 2022; Park et al., 2023b; Rho et al., 2024), we sampled a skill latent at the beginning of each episode and kept it fixed throughout the rollout. While transitioning between skills is orthogonal and can be incorporated into both our method and the baselines to further improve robustness, we excluded it here to focus on a clean understanding of algorithmic contributions.
@@ G.4 Hyperparameters
Optimizer Adam(Kingma and Ba, 2014)